OpenStudy (anonymous):

m^2- 1 7 ----m = ---- 2 2 Solve by completing the square

jhonyy9 (jhonyy9):

this is m*2 -m/2 =7/2 ?

jhonyy9 (jhonyy9):

so if yes we have 2m*2 -m -7=0

jhonyy9 (jhonyy9):

1 +/- radical (1 +56) 1+/- radical 57 m_1,_2 = -------------------- = --------------- 4 4

OpenStudy (anonymous):

so that is what I need to put as the answer then?

OpenStudy (anonymous):

hello is that what I need then....????

OpenStudy (anonymous):

I am so confused when you write it this way?

OpenStudy (anonymous):

can you answer me please jhonyy9?

jhonyy9 (jhonyy9):

ok sorry yes those are solutions